### v1

# Localized Triplons and Site Stuffing in the Quantum Dimer Magnet BiYbGeO5

Abstract
Thermodynamic and muon spin-relaxation measurements have recently highlighted BiYbGeO5 as a new example of a rare-earth-based quantum dimer magnet with isolated Yb3+ spin- dimers. However, direct spectroscopic evidence of the triplet excitations and measurements of the structural disorder are lacking. In this work, polycrystalline BiYbGeO5 was synthesized using conventional high-temperature solid-state methods and investigated via high-resolution neutron powder diffraction and inelastic neutron scattering. Diffraction measurements down to 58 mK reveal no signatures of magnetic order and indicate that nearly 20% of Yb3+ sites are replaced by non-magnetic Bi3+, introducing significant structural disorder. Inelastic neutron scattering shows dispersionless triplon excitations, consistent with localized, non-interacting spin dimers. Fits to the triplet excitation spectrum identify an XXZ-type anisotropic exchange with = 0.11(2) meV and meV. These findings establish BiYbGeO5 as a structurally disordered but magnetically well-isolated quantum dimer system, providing a model platform for studying the resilience of entangled spin states to site dilution.
I Introduction
Many exotic quantum states in condensed matter systems, such as quantum spin liquids, emerge from the entanglement of electron pairs into spin singlets. In a quantum spin liquid, the ground state can be described as a superposition of all possible singlet coverings of the lattice Balents (2010). Understanding how singlets form and respond to perturbations is therefore essential for identifying and designing new quantum spin liquid candidates.
In cases where long-range entanglement is absent but local singlets form between neighboring spins without breaking lattice symmetries, the resulting state is known as a quantum dimer magnet. These systems are characterized by the absence of long-range magnetic order down to the lowest temperatures and the presence of coherent spin excitations – called triplons. In idealized models with negligible interdimer interactions, the singlet-triplet gap can be closed by an external magnetic field, leading directly to a fully polarized paramagnet. However, real materials invariably host finite interdimer interactions that can stabilize an intermediate field-induced phase, which maps onto a Bose-Einstein condensate of triplons Zapf et al. (2014). Quantum dimer magnets thus serve as a minimal platform for exploring quantum entanglement and the onset of quantum criticality in low-dimensional spin systems.
To date, most quantum dimer magnets have been discovered in 3 transition metal-based compounds, where strong exchange interactions often result in prohibitively high magnetic field strengths (> 20 T) to close the single-triplet gap. The discovery of Yb2Si2O7 in 2019 Hester et al. (2019) marked the first rare-earth-based example of such a state, where the contracted nature of the 4 orbitals leads to a relatively small magnetic field to close the singlet-triplet gap (1.0 T). Since the discovery of Yb2Si2O7 several other 4-based quantum magnets with emergent singlet ground states have been proposed in Yb2SiO5 Hase et al. (2025), Yb-doped YAlO3 Nikitin et al. (2020), Yb2Be2SiO7 Brassington et al. (2025), and BiYbGeO5 Mohanty et al. (2023).
BiYbGeO5 is of particular interest as magnetometry, heat capacity, and muon spin relaxation measurements suggest the formation of isolated spin singlets composed of effective spin-1/2 Yb3+ moments Mohanty et al. (2023). Field-dependent heat capacity measurements by Mohanty et al. indicate anisotropic intradimer exchange with 1.3 K and J 2.6 K with no thermodynamic evidence for interdimer interactions. These measurements confirm the absence of magnetic order at low temperatures and the presence of ground state singlets; however, the ideal method for confirming the presence of triplon excitations – and their energy scale – is inelastic neutron scattering, which has not been undertaken until this study.
In this work, we investigate the proposed dimer ground state in BiYbGeO5 using inelastic neutron scattering to search for coherent triplon excitations and assess the degree of Bi3+/Yb3+ site disorder through high-resolution neutron diffraction. These measurements allow us to probe the strength of the intradimer and interdimer exchange and investigate how the triplon excitations survive in the presence of disorder, a key question for understanding the robustness of dimer physics beyond the clean limit.
II Experimental Methods
II.1 Synthesis
A pale-yellow polycrystalline powder of BiYbGeO5 was synthesized using the solid-state synthesis method. This method is an adaptation from the synthesis method proposed by Cascales et al. Cascales et al. (2002). Bi2O3 (99.99%, Thermo Scientific), Yb2O3 (99.9%, Sigma-Aldrich), and GeO2 (99.99%, Sigma-Aldrich) were dried to remove any water of crystallization, after which they were stored in a desiccator. Stoichiometric amounts of the dried reactants were taken in molar ratios of in an agate mortar and pestle and ground, pressed into pellets, heated in air in an alumina crucible inside a box furnace for 24 hours each, and quenched 111It was found with further synthesis that quenching is not necessary and the pellets can be cooled at a slower rate as well. This process was repeated four times, where the temperature was set to C, C, C, and C for each heating, respectively. For the last two heatings at C, the pellets were pressed by adding 1 ml of PVA (polyvinyl alcohol) solution (50 mg/ml) to facilitate proper pellet formation. Phase purity was assessed using powder X-ray diffraction (XRD) at room temperature using a Rigaku SmartLab X-ray Diffractometer (CuKα source with Kα1/Kα2 of 1.540510Å/1.544330Å).
II.2 Neutron Powder Diffraction
Neutron diffraction measurements were performed on powdered BiYbGeO5 using the HB-2A Neutron Powder Diffractometer (POWDER) at the High Flux Isotope Reactor at Oak Ridge National Laboratory and analyzed using the Rietveld refinement method within the FullProf Suite Rodríguez-Carvajal (1993). Approximately 5 g of the compound was loaded in a copper canister backfilled with 10 atm of He to ensure proper thermalization at low temperatures. Data was collected at 20 K, 0.9 K, and 0.058 K. Measurements were taken using two incident wavelengths of = 2.41 Å and = 1.54 Å using the Ge(113) and Ge (115) monochromators, respectively. For all measurements, the collimator setup was open-open-12’. A measurement with = 1.54 Å provided a larger Q range, allowing for more precise measurement of potential site mixing between Bi3+ and Yb3+.
II.3 Inelastic Neutron Scattering
Inelastic neutron scattering (INS) spectra were measured at the Spallation Neutron Source at Oak Ridge National Laboratory using the BL-14B Hybrid Spectrometer (HYSPEC). Slices and cuts of the data were made using the MANTID analysis software Arnold et al. (2014). Data was collected at low temperature (0.250 K) to observe triplet excitations and at high temperature (100 K) as a background measurement. Multiple datasets were also collected at intermediary temperatures (0.580 K, 0.850 K, 1.5 K, 2.5 K, 10 K, and 20 K) to analyze the evolution of the triplet excitation intensity with increasing temperature. All measurements were performed under no external magnetic field in the unpolarized mode. The Fermi chopper speed was set at 420 Hz, and the lowest incident neutron energy of 3.8 meV was chosen to give the best possible resolution of 0.1 meV at the elastic line. All data, except for 0.250 K, were collected using a detector bank covering a scattering angle (2) range of , giving horizontal detector coverage of 0.20 Å-1 to 1.5 Å-1 at the elastic line. To better measure the structure factor of the dimers at 250 mK, an extended detector coverage of was used to give a |Q| range of 0.20 Å-1 to 2.0 Å-1 at the elastic line.
III Results
III.1 Neutron Powder Diffraction
Refinement of the 20 K dataset yielded lattice parameters of a = 5.2866(1) Å, b = 15.1762(3) Å, and c = 10.9333(2) Å, smaller than previously reported room-temperature values by Mohanty et al. Mohanty et al. (2023) and Cascales et al. Cascales et al. (2002). This contraction of the unit cell upon cooling is consistent with typical thermal contraction behavior.
Structural refinements of the = 1.54 Å scan at 20 K, further revealed a degree of site disorder between the Bi3+ and Yb3+ ions, likely driven by the similar ionic radii of Bi3+ (1.17 Å) and Yb3+ (0.985 Å) Shannon (1976), their identical valence states, and that all atoms occupy the same Wyckoff position, imposing identical symmetry constraints. Approximately 19.2(26)% of the Yb3+ (occupying the Yb site, see Table 1) are found to be replaced by Bi3+, as determined by Rietveld analysis. Alternative scenarios, including Yb3+ occupancy of Bi3+ sites and bidirectional site mixing, were tested but yielded unrealistic parameters. Given that Bi3+ is non-magnetic, this substitution effectively breaks Yb-Yb magnetic dimers, generating an equivalent fraction of unpaired Yb3+ spins. Such site mixing defects are expected to locally increase the average interdimer distances and may play a crucial role in the low-temperature magnetic behavior of BiYbGeO5.
The absence of additional Bragg reflections across the entire measured temperature range confirms that all peaks are structural in origin, indicating a lack of long-range magnetic order down to 58 mK. To probe the possibility of short-range magnetic correlations, the diffraction pattern measured at 20 K was subtracted from the 0.058 K pattern. The resulting difference was featureless within the noise level, indicating the absence of any diffuse scattering and ruling out the presence of incipient magnetic transitions or short-range order.
Two temperature-independent impurity peaks are visible at Å-1 and Å-1 with very small intensities. These peaks are not visible in the shorter-wavelength ( = 1.54 Å) scan shown in Figure 2, likely due to reduced resolution at higher |Q| values where peak overlap becomes more significant. The impurity peaks are visible in the longer wavelength neutron scans and the XRD scan, which have been provided in the supplemental information SI .
Atom Site x y z B Occupancy Bi1 8c 0.95685(69) 0.23741(24) 0.14649(38) 0.130(66) 1 Yb 8c 0.00808(65) 0.05396(15) 0.35944(30) 0.042(50) 0.808(26) Bi2 8c 0.00808(65) 0.05396(15) 0.35944(30) 0.042(50) 0.192(26) Ge 8c 0.00543(93) 0.40456(27) 0.40269(34) 0.282(62) 1 O1 8c 0.07054(97) 0.30183(32) 0.32959(48) 0.21(11) 1 O2 8c 0.3002(11) 0.43740(39) 0.45670(53) 0.19(8) 1 O3 8c 0.3182(14) 0.12552(40) 0.47399(56) 0.62(13) 1 O4 8c 0.2542(13) 0.15041(43) 0.23658(57) 0.50(10) 1 O5 8c 0.3493(11) 0.47814(43) 0.19261(54) 0.37(10) 1
III.2 Inelastic Neutron Scattering
To investigate the low-energy spin dynamics in BiYbGeO5, INS measurements were carried out using HYSPEC. As shown in Figures 3a-c, a well-defined excitation band is observed across a broad range of momentum transfers, |Q|. This excitation appears dispersionless within the instrumental resolution of 0.1 meV, consistent with isolated dimers with negligible interdimer exchange. Energy transfers were measured up to 3.61 meV (see supplemental information SI ), in which no crystalline electric field (CEF) excitations were detected, placing a lower bound on the CEF gap of 42 K.
The scattering intensity was plotted as a function of energy transfer (E) by integrating a |Q| range of 0.25 Å-1 – 1.4 Å-1 as shown in Figure 3d. Although the triplon excitation energy remains temperature-independent, its spectral weight decreases with increasing temperature, consistent with thermal depopulation of the singlet ground state.
To isolate the triplon excitation, the high-temperature (100 K) data were subtracted from the low-temperature (0.250 K) dataset, effectively removing the elastic background and sample environment. The resulting difference spectra show a single excitation that is limited by the instrument resolution. A two-Gaussian model was considered to estimate the intradimer exchange interactions, consistent with the expectations of an XXZ-type anisotropic exchange in isolated dimers. This model provided the best agreement with the INS data, as shown in Figure 4. Single and triple-Gaussian models were also tested, leading to poor and unrealistic fits, respectively. The single and triple-Gaussian fits, with their fitted parameters, have been included in the SI SI . In order to more precisely determine the exchange interactions, a higher-resolution INS scattering experiment would be required.
In the XXZ dimer framework at zero magnetic field, three triplet excitations are expected: a single state and a doubly degenerate excitation to the states, with the latter state expected to have a larger spectral weight. Under this assumption, the extracted exchange parameters were found to be meV and meV, yielding an anisotropy ratio of . The energy hierarchy agrees qualitatively with that proposed by Mohanty et al. Mohanty et al. (2023), which was inferred from heat capacity measurements. However, while the values are comparable within experimental uncertainty, the extracted from inelastic neutron scattering is significantly lower than previously reported. This discrepancy emphasizes the utility of direct spectroscopic probes such as INS for resolving anisotropic exchange strengths and may reflect inherent limitations of thermodynamic fitting approaches.
Exchange interactions This Work (meV) Mohanty et al. (meV) 0.11(2) 0.11 0.15(1) 0.22
The -dependence of the triplet excitation at 0.250 K was further analyzed to confirm the spatial configuration of Yb dimers as shown in Figure 5. The measured intensity was fit to the magnetic structure factor for an isolated dimer,
where, is the magnetic form factor for Yb3+ and is the intradimer distance. The magnetic form factor was modeled as Lovesey and Lynn (1986), using tabulated coefficients for Yb3+ Freeman and Desclaux (1979). The extracted dimer distance of 3.499(11) Å is in close agreement with the nearest-neighbor Yb-Yb spacing obtained from the Rietveld refinement of the neutron diffraction data collected at 58 mK (3.495(9) Å), confirming that the observed magnetic excitations originate from nearest-neighbor Yb dimers.
IV Discussion
BiYbGeO5 exhibits signatures of an isolated quantum dimer magnet, with a significant degree of Bi–Yb site mixing ( 19.2%) leading to broken Yb-Yb dimers. The substitution of non-magnetic Bi3+ for Yb3+ both increases the average separation between remaining magnetic dimers and may play a role in further suppressing interdimer triplon hopping, accounting for the flat band observed in inelastic neutron scattering. This level of site mixing occurs despite stoichiometric synthesis, implying that excess Bi3+ must occupy Yb3+ sites and expelled Yb3+ segregates into minor impurity phases. The number and intensity of these impurity peaks are too low to enable precise identification of the impurity phases. No signs of magnetic excitations from the impurity phase are observed in the inelastic neutron scattering data obtained from HYSPEC, likely due to the relatively small amount of impurity.
Previous work by Mohanty et al. Mohanty et al. (2023) and Cascales et. al Cascales et al. (2002) utilized X-ray diffraction data of BiYbGeO5 to assess phase purity and therefore they do not report any Yb/Bi site mixing. However, the x-ray cross-sections of Bi3+ ( barns) and Yb3+ ( barns) are fairly close, as that is predominantly a function of the number of electrons in the ions Seltzer (1987). The neutron cross-sections, by contrast, are quite distinct (9.3148 barns for Bi and 19.42 barns for Yb), allowing one to discern the site mixing happening between the two ions Varley (1992). The presence of site mixing is perhaps unsurprising as Cascales et al. also analyzed the XRD pattern of BiYGeO5, which showed 12% of Bi3+ stuffing in the Y3+ site. In the work of Mohanty et al. they report an impurity contribution of that is required to appropriately fit their field-dependent heat capacity data, assuming the impurity spin to be . The presence of this impurity may be explained by Bi3+ stuffing into the Yb3+ site where the substitution of a Yb-Yb dimer by a Bi-Yb pair results in an unpaired Yb3+ spin, which could yield this measured "impurity" contribution to magnetic susceptibility.
Site mixing effects of this sort are well known in Yb-based magnets and can dramatically alter magnetic behavior. For example, the pyrochlore quantum spin ice Yb2Ti2O7 is notoriously sensitive to small off-stoichiometry. Even 2% “stuffing” of the Ti4+ sites by extra Yb3+ leads to strongly sample-dependent ground states Gaudet et al. (2015); Ghosh and Manousakis (2018). Stoichiometric polycrystalline Yb2Ti2O7 shows a sharp transition, whereas lightly stuffed single crystals exhibit broad heat-capacity anomalies and no conventional order Ross et al. (2012); demonstrating that magnetic properties can differ significantly between single crystals and polycrystalline samples due to subtle synthesis-dependent variations in disorder. By analogy, BiYbGeO5’s site mixing similarly introduces random disruption, but in this case, the disorder further suppresses interdimer triplon hopping rather than creating competing couplings. Future growth of BiYbGeO5 single crystals will be crucial to determine whether this site mixing is an intrinsic feature of the material or a consequence of polycrystalline synthesis.
Another instructive comparison is the Shastry–Sutherland, isolated dimer system Yb2Be2SiO7. In this system, the site mixing between the non-magnetic Be2+ and Si4+ ions leads to a significant broadening of the triplon band as observed in INS measurements Brassington et al. (2025). Er2Be2SiO7, by contrast, does not show the Be/Si site mixing and develops long-range magnetic ordering upon cooling Brassington et al. (2024). Although the disorder in Yb2Be2SiO7 is not at the magnetic site, it still perturbs the effective exchange landscape.
Beyond the experimental implications, the isolated dimers of BiYbGeO5 may be an ideal system for testing quantum simulations of entangled states in condensed matter systems. Quantum dimer magnets serve as an ideal two-qubit prototype for quantum simulation, directly mapping each spin in a dimer to a qubit pair and encoding the antiferromagnetic exchange interaction into entangling gates Ardavan et al. (2015); Jing-Min et al. (2005). Eassa et al. demonstrated that resource-efficient fast-forwarding measurement performed with short-depth circuits can faithfully reproduce the dynamical structure factor and the triplon gap of dimers on current quantum processors, with fidelities high enough to match costly INS measurements Eassa et al. (2024). BiYbGeO5 may be an ideal system for extending that analysis to rare-earth-based quantum magnets.
V Conclusions
In this work, polycrystalline samples of BiYbGeO5 were synthesized via conventional solid-state reactions, and their magnetic and structural properties were investigated through neutron powder diffraction and inelastic neutron scattering. Neutron diffraction measurements carried out at 20 K, 0.9 K, and 58 mK, confirming the absence of magnetic orderings. Structural refinements revealed significant site disorder, with 19.2(26)% of Yb3+ sites replaced by non-magnetic Bi3+ ions. Inelastic neutron scattering measurements uncovered dispersionless triplon excitations, consistent with localized triplet modes due to isolated Yb–Yb dimers. Gaussian fits to the energy spectrum yielded an XXZ-type anisotropic exchange interaction, with meV and meV. Analysis of the magnetic structure factor further determined the intradimer Yb-Yb distance to be 3.499(11) Å, in excellent agreement with the nearest-neighbor spacing extracted from diffraction. These results confirm that despite substantial structural disorder, BiYbGeO5 hosts a magnetically disordered quantum ground state composed of well-isolated Yb dimers, making it a compelling platform for exploring disorder-resilient quantum magnetism.
VI Acknowledgements
R. Kapoor and G. Hester would like to thank R. Ganesh for insightful conversations closely and distantly related to this work. We would also like to thank C. Sarkis for facilitating the measurements using HYSPEC for this work. We also extend our gratitude to Carter Fortuna for synthesizing many variants of this material. A portion of this work used resources at the High-Flux Isotope Reactor and the Spallation Neutron Source, which are DOE Office of Science User Facilities operated by Oak Ridge National Laboratory. The beamtime was allocated to POWDER (HB-2A) and HYSPEC (BL-14B) on proposal numbers IPTS-32979 and IPTS-35019, respectively.